Circular No. 6761

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD P. Garnavich, R. Kirshner, and P. Challis, Harvard-Smithsonian Cen-

ter for Astrophysics, report on behalf of the Supernova Intensive Study team: “Long-slit spectra of SN 1987A and its circumstellar emission rings were obtained with the Space Telescope Imaging Spectrograph aboard the Hubble Space Telescope (HST). Ultraviolet spectra (range 113–170 nm) were obtained on Sept. 27 UT, and low-resolution optical spectra (range 530–1000 nm) were taken on Oct. 6. The spectra show evidence of both the interaction of the supernova ejecta with the inner H II region and the collision between the ejecta and the circumstellar ring. The 0′′.5-wide slit for the ultraviolet spectra was set at p.a. 33^{o} to include the ‘hot spot’ (cf. IAUC 6665, 6710), which is a protrusion of the inner ring being shocked by the supernova ejecta. Emission lines of N V, O IV, N IV, C IV, and He II are detected at the hot-spot location 0′′.1 interior to the ring. The presence of C IV emission, which is normally absorbed by interstellar gas, indicates the emission is blueshifted relative to the rest of the ring. Bright Lyα emission is seen as a blueshifted arc on the north side of the ring and a fainter, redshifted arc on the south side. At low velocity, the blueshifted arc extends to within 0′′.1 of the ring and reaches peak brightness at 11 000 km/s with a projected distance of 0′′.2 from the supernova position. The blueshifted arc becomes undetectable beyond 15 000 km/s at the position of the supernova. The brightness difference between the redshifted and blueshifted Lyα emission implies either an asymmetry in the interaction between the near side and far side, or an increase in brightness on a timescale faster than the light travel time across the ring. Weak emission redward of Lyα is possibly due to N V 124-nm, as predicted by Michael et al. (1997, Ap.J., submitted). The HST optical spectrum taken with a 0′′.2 slit shows a highvelocity, blueshifted arc of Hα, corresponding to the Lyα emission. The Hα arc begins 0′′.1 inside of the peak of the ring flux and is clearly visible to 12 000 km/s. The redshifted arc is significantly fainter and visible only to 5000 km/s.”

COMET C/1997 T1 (UTSUNOMIYA) H. Kawakita, Osaka, Japan; and M. Fujii, Okayama, Japan, obtained

a low-dispersion (resolution ∼ 3 nm, range 380–730 nm) spectrogram of this comet with the 1.01-m telescope at Bisei Astronomical Observatory on Oct. 11.6 UT. CN, C_{3}, C_{2}, NH_{2} and [O I] emissions are present.

1997 October 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green